# Optimized matching conditions for self-guided laser wakefield accelerators

### v1

@@ 1 Introduction
Laser wakefield acceleration (LWFA) has emerged as a rapidly advancing technique capable of achieving acceleration gradients several orders of magnitude higher than those attainable with conventional radio-frequency accelerators [1, 2, 3, 4]. The underlying mechanism relies on the interaction between an intense laser pulse and a plasma, where the laser drives a large-amplitude plasma wave that can trap and accelerate electrons to relativistic energies. Over recent decades, LWFA has attracted considerable attention owing to its potential to enable compact, high-energy electron accelerators with broad applications, including advanced radiation sources [5, 6, 7, 8, 9], muon sources [10, 11, 12, 13], radiotherapy [14, 15, 16, 17, 18], and the exploration of strong-field quantum electrodynamics phenomena [19, 20, 21, 22].
@@
To address these challenges, research has increasingly turned to advanced computational and optimization techniques. Modern high-performance computing enables the modeling of laser–plasma interactions that capture the essential physical phenomena with high fidelity and reasonable computational cost. However, the number of simulations required to explore the high-dimensional and wide-ranging parameter space remains beyond current capabilities. To overcome this limitation, artificial intelligence and machine learning methods are playing an increasingly important role in efficiently navigating the parameter space, identifying optimal operating regimes, and reducing computational expense [23, 24, 25, 26, 27, 28].
In the present work, we revisit the so-called matching conditions for self-guided laser pulse propagation in plasma [29, 30, 31, 32] and examine whether their formulation can be refined to enhance the maximum energy attainable by electrons accelerated through LWFA. At this stage, we restrict our analysis to the optimization of maximum electron energy alone, disregarding other beam parameters. This focus is motivated by specific applications in which exceeding a threshold energy is essential, such as muon production [10, 11, 12, 13] and nuclear activation [33, 34].
To accomplish our goal, we employ Bayesian optimization (BO), a machine learning method particularly well suited for optimizing expensive-to-evaluate objective functions (i.e., functions to be minimized or maximized). In our study, the objective function—the maximum electron energy—is evaluated through a series of particle-in-cell (PIC) simulations using a quasi-three-dimensional (quasi-3D) geometry [35, 36] and a Lorentz-boosted frame [37, 38, 39]. We note that BO has already been successfully applied to complex optimization problems in LWFA [40, 41, 42, 43, 44, 45, 46, 47, 48], and that both the quasi-3D geometry and the Lorentz-boosted-frame simulation techniques have been well established in previous studies [49, 50, 39, 11, 51].
@@ 2 Description of Gaussian laser pulse
Here, $r_{\mathrm{e}}=e^{2}/4\pi\epsilon_{0}m_{\mathrm{e}}c^{2}$ denotes the classical electron radius, with $\epsilon_{0}$ being the vacuum permittivity. The quantities in Eq. (5) evaluate to $\overline{\mathcal{E}}\approx 29~\mathrm{\upmu J}$ (for $\lambda_{0}=1~\mathrm{\upmu m}$) and $\overline{\mathcal{P}}\approx 17.4~\mathrm{GW}$. The corresponding laser power is then expressed as
@@ 3 Self-guiding and matching conditions
For efficient LWFA, the laser pulse must preserve high intensity in plasma over distances significantly longer than those limited by diffraction. To counteract diffraction, guiding of the laser pulse within the plasma is essential. In general, this can be achieved either through external optical guiding or through self-guiding mechanisms. In the present work, we focus on the latter case. Self-guiding is experimentally attractive because it does not require precise control of external guiding structures, such as capillary-discharge waveguides [52, 53, 54, 55, 56, 57, 3] or hydrodynamically formed plasma channels [58, 59, 60, 61, 62, 4]. On the other hand, it typically operates in a strongly nonlinear regime, which can lead to a reduction in the quality and stability of the accelerated electron beam compared with externally guided configurations [2, 4]. These factors make self-guided LWFA a promising and relatively straightforward target for optimization.
@@ 4 Generalized matching conditions and parameterization
To reduce the number of free parameters involved in the optimization process, we adopt the following simplifying assumptions: (i) LWFA is driven by a linearly polarized Gaussian laser pulse with a constant spectral phase, specified central wavelength, and total energy; (ii) the plasma has a uniform density profile; (iii) the acceleration operates in the matched regime defined by Eqs. (11) and (12); and (iv) electrons are injected into the accelerating structure externally. These assumptions allow us to isolate the essential physical dependencies while minimizing computational complexity. It should be noted, however, that under these constraints the maximum achievable electron energy is not necessarily optimal. More complex configurations, such as LWFA driven by non-Gaussian [70, 71] or frequency-chirped [72, 73] laser pulses, tailored plasma profiles [74, 75, 76, 77], or operation in a mismatched regime [78, 79], can potentially yield higher energy gains, albeit at the cost of introducing additional degrees of freedom and complicating the optimization process. In future work, the methods developed in the present study may be extended to include these more general cases.
@@
In this study, we select a laser pulse energy of $10~\mathrm{mJ}$ and a wavelength of $1~\mathrm{\upmu m}$. Although the chosen energy is relatively low for typical LWFA experiments, it ensures that the corresponding acceleration length remains moderate (less than $200~\mathrm{\upmu m}$ in all simulated cases), thereby keeping the computational cost of the PIC simulations used for BO tractable. On the other hand, such low-energy pulses correspond to ultrashort durations, for which carrier–envelope phase effects can influence the wakefield dynamics and electron acceleration [80, 81, 82, 83]. The impact of these effects and the corresponding scalability of the optimization to high-energy, long-pulse regimes will be examined in future work.
@@ 5 Particle-in-cell simulations
To isolate the acceleration dynamics from the specifics of various injection mechanisms, an external beam of test electrons is introduced. The test electrons move self-consistently according to their charge and mass but do not contribute any current to the plasma, effectively acting as passive tracers. This approach is valid for regimes well below the beam-loading threshold [84, 85], where the influence of the accelerated electrons on the wakefield structure is negligible. The initial velocity of the test electrons is set to $v_{0}=c\sqrt{1-\omega_{\mathrm{p}}^{2}/\omega_{0}^{2}}$, approximately matching the phase velocity of the plasma wave to ensure efficient electron trapping. The transverse and longitudinal dimensions of the electron beam are chosen such that it fully overlaps with the generated wakefield structure.
@@ 6 Bayesian optimization
In our setup, a Gaussian process (GP) [90] serves as the surrogate model, following standard practice in BO. The GP relies on a kernel (or covariance function) that defines correlations between points in the parameter space and encodes prior assumptions about the smoothness and general behavior of the objective function. We adopt the radial basis function kernel [91], which assumes the objective function to be infinitely differentiable—an appropriate choice given the expected smooth dependence of the output on the input parameters.
The acquisition function employed is the Monte Carlo–based batched upper confidence bound ($q$-UCB) [92, 93]. The $q$-UCB acquisition function includes a tunable hyperparameter $\beta_{q\text{-}\mathrm{UCB}}$, which governs the balance between exploration and exploitation. We set $\beta_{q\text{-}\mathrm{UCB}}=10$ and keep it constant throughout the optimization process, as this value has proven to provide a good balance in a series of preliminary tests. Furthermore, we employ a batched acquisition strategy that allows multiple trials to be selected and evaluated concurrently.
@@
Based on our previous work [48], the optimum is expected to lie within these intervals; hence, the parameter limits are not dynamically adjusted during the optimization process. The single optimization objective is the maximum test-electron macroparticle energy achieved in the simulation. Additionally, the acceleration length—defined as the propagation distance in plasma at which the maximum energy is reached—is recorded for post-analysis, although it is not included in the optimization objective.
@@ 7 Optimization results
This corresponds to the following set of initial laser and plasma parameters: $a_{0}\approx 2.3$, $\tau_{0}\approx 9.5~\mathrm{fs}$ (yielding $\mathcal{P}_{0}\approx 1~\mathrm{TW}$), $w_{0}\approx 3~\mathrm{\upmu m}$, and $n_{\mathrm{e}}\approx 3\times 10^{19}~\mathrm{cm^{-3}}$. Across the explored parameter space, the maximum electron energy ranges from $\approx 25$ to $77~\mathrm{MeV}$, with the corresponding acceleration distance varying between $\approx 19$ and $192~\mathrm{\upmu m}$.
@@
This indicates that near-maximum electron energies can be produced across several different sets of initial parameters. For instance, in the present $10~\mathrm{mJ}$ case, electrons in the top $5\%$ of the energy distribution are obtained either for $\tau_{0}\approx 9.8~\mathrm{fs}$, $w_{0}\approx 2.4~\mathrm{\upmu m}$, and $n_{\mathrm{e}}\approx 3.4\times 10^{19}~\mathrm{cm^{-3}}$, or for $\tau_{0}\approx 8.1~\mathrm{fs}$, $w_{0}\approx 3.3~\mathrm{\upmu m}$, and $n_{\mathrm{e}}\approx 2.9\times 10^{19}~\mathrm{cm^{-3}}$. Furthermore, by fixing the pulse duration and plasma density near their optimal values (i.e., $\tau_{0}\approx 9.5~\mathrm{fs}$ and $n_{\mathrm{e}}\approx 3\times 10^{19}~\mathrm{cm^{-3}}$), variations of the focal spot from $\approx 2.4$ to $3.3~\mathrm{\upmu m}$ still yield near-maximum electron energies. These findings demonstrate that operation within the optimal LWFA regime (i) can be realized through multiple parameter combinations, providing substantial flexibility for experimental implementation, and (ii) can be maintained without the need for precisely tuned laser or plasma parameters, thereby significantly relaxing the operational constraints.
@@
The simulation results indicate that achieving efficient self-guiding under realistic conditions requires accounting for the self-consistent evolution of the laser pulse and the plasma. As discussed in Ref. [48], this implies that (i) the initial laser power should be set slightly above the critical power and (ii) the pulse duration be somewhat longer than the resonant duration for linear plasma wave excitation. We note that, for a Gaussian pulse with $a_{0}\ll 1$, the resonant duration can be calculated analytically as $\tau_{r}\omega_{\mathrm{p}}\approx 2\sqrt{2\ln{2}}\approx 2.35$ [96].
In this work, we build upon our previous study [48], where BO was applied to the case of $\kappa=2$, the electron injection was realized using the nanoparticle-assisted method [97], and all other laser and plasma parameters were kept identical. The adoption of advanced numerical techniques in the present analysis enabled a more thorough exploration of the parameter space, resulting in a maximum electron energy almost 15% higher than previously reported.
@@ 8 Conclusion
The optimization revealed that a self-guided LWFA driven by a $10~\mathrm{mJ}$ laser pulse in a uniform plasma is, in principle, capable of producing electrons with energies approaching $80~\mathrm{MeV}$ over an acceleration distance below $200~\mathrm{\upmu m}$. The maximum energy was achieved for a proportionality parameter close to the widely accepted factor of 22 ($\pm 0.4$). Furthermore, electrons with energies near the maximum value were obtained across a relatively broad range of input parameters and without the need for their precise tuning. For instance, when the pulse duration and plasma density are set to $9.5~\mathrm{fs}$ and $3\times 10^{19}~\mathrm{cm^{-3}}$, respectively, varying the focal spot between $2.4$ and $3.3~\mathrm{\upmu m}$ still yields near-maximum electron energies. This provides substantial flexibility for experimental implementation and significantly relaxes the operational constraints associated with self-guided LWFA. Although the optimization is demonstrated for a relatively low-energy driver pulse, the parameter space is defined in terms of dimensionless quantities, for which the location and extent of the optimal operating region are expected to depend only weakly on the laser energy due to the self-similar nature of the underlying physics.

### v2

@@ 1 Introduction
Laser wakefield acceleration (LWFA) has emerged as a rapidly advancing technique capable of achieving acceleration gradients several orders of magnitude higher than those attainable with conventional radio-frequency accelerators [1]. The underlying mechanism relies on the interaction between an intense laser pulse and a plasma, where the laser drives a large-amplitude plasma wave that can trap and accelerate electrons to relativistic energies [2]. Over recent decades, LWFA has attracted considerable attention owing to its potential to enable compact, high-energy electron accelerators [3, 4] with broad applications, including advanced radiation sources [5, 6], muon sources [7, 8, 9], radiotherapy [10, 11, 12], and the exploration of strong-field quantum electrodynamics phenomena [13, 14].
@@
To address these challenges, research has increasingly turned to advanced computational and optimization techniques. Modern high-performance computing enables the modeling of laser–plasma interactions that capture the essential physical phenomena with high fidelity and reasonable computational cost. However, the number of simulations required to explore the high-dimensional and wide-ranging parameter space remains beyond current capabilities. To overcome this limitation, artificial intelligence and machine learning methods are playing an increasingly important role in efficiently navigating the parameter space, identifying optimal operating regimes, and reducing the total computational expense [15, 16].
In the present work, we revisit the so-called matching conditions for self-guided laser pulse propagation in plasma [17, 18, 19, 20] and examine whether their formulation can be refined to enhance the maximum energy attainable by electrons accelerated through LWFA. At this stage, we restrict our analysis to the optimization of maximum electron energy alone, disregarding other beam parameters. This focus is motivated by specific applications in which exceeding a threshold energy is essential, such as muon production [7, 8, 9] and nuclear activation [21, 22]. It is important to note, however, that the electron energy is expected to depend predominantly on the parameters of laser and the background plasma, whereas beam quality metrics (e.g., charge, energy spread, and divergence) are more strongly influenced by the specific electron injection mechanism. Consequently, the optimal conditions for maximizing electron energy identified here are expected to remain valid across different controllable injection schemes, while optimization of beam quality for a given injection mechanism can be performed subsequently.
To accomplish our goal, we employ Bayesian optimization (BO), a machine learning method particularly well suited for optimizing expensive-to-evaluate objective functions (i.e., functions to be minimized or maximized). In our study, the objective function—the maximum electron energy—is evaluated through a series of particle-in-cell (PIC) simulations using a quasi-three-dimensional (quasi-3D) geometry [23, 24] and a Lorentz-boosted frame [25, 26, 27]. We note that BO has already been successfully applied to complex optimization problems in LWFA [28, 29, 30, 31, 32], and that both the quasi-3D geometry and the Lorentz-boosted-frame simulation techniques have been well established in previous studies [33, 34, 27, 35].
@@ 2 Description of Gaussian laser pulse
Here, $r_{\mathrm{e}}=e^{2}/4\pi\epsilon_{0}m_{\mathrm{e}}c^{2}$ denotes the classical electron radius, with $\epsilon_{0}$ being the vacuum permittivity. The quantities in Eq. (5) evaluate to $\overline{\mathcal{E}}\approx 29~\mathrm{\upmu J}$ (for $\lambda_{0}=1~\mathrm{\upmu m}$) and $\overline{\mathcal{P}}\approx 17.4~\mathrm{GW}$. The corresponding laser power is then expressed as [36]
@@ 3 Self-guiding and matching conditions
For efficient LWFA, the laser pulse must preserve high intensity in plasma over distances significantly longer than those limited by diffraction. To counteract diffraction, guiding of the laser pulse within the plasma is essential. In general, this can be achieved either through external optical guiding or through self-guiding mechanisms. In the present work, we focus on the latter case. Self-guiding is experimentally attractive because it does not require precise control of external guiding structures, such as capillary-discharge waveguides [37, 38] or hydrodynamically formed plasma channels [39, 40]. On the other hand, it typically operates in a strongly nonlinear regime, which can lead to a reduction in the quality and stability of the accelerated electron beam compared with externally guided configurations [2, 4]. These factors make self-guided LWFA a promising and relatively straightforward target for optimization.
@@ 4 Generalized matching conditions and parameterization
To reduce the number of free parameters involved in the optimization process, we adopt the following simplifying assumptions: (i) LWFA is driven by a linearly polarized Gaussian laser pulse with a constant spectral phase, a specified central wavelength, and total energy; (ii) the plasma has a uniform density profile; (iii) the acceleration operates in the matched regime defined by Eqs. (11) and (12); and (iv) accelerated electrons are treated as test particles and are injected into the accelerating structure externally. These assumptions allow us to isolate the essential physical dependencies while minimizing computational complexity. It should be noted, however, that under these constraints the maximum achievable electron energy is not necessarily optimal. More complex configurations, such as LWFA driven by non-Gaussian [48, 49] or frequency-chirped [50, 51] laser pulses, tailored plasma profiles [52, 53], or operation in a mismatched regime [54, 55], can potentially yield higher energy gains, albeit at the cost of introducing additional degrees of freedom and complicating the optimization process. In future work, the methods developed in the present study may be extended to include these more general cases.
@@
In this study, we select a laser pulse energy of $10~\mathrm{mJ}$ and a wavelength of $1~\mathrm{\upmu m}$. Although the chosen energy is relatively low for typical LWFA experiments, it ensures that the corresponding acceleration length remains moderate (less than $200~\mathrm{\upmu m}$ in all simulated cases), thereby keeping the computational cost of the PIC simulations used for BO tractable. Furthermore, the selected laser parameters closely match the performance of state-of-the-art ultrafast fiber laser technology [56], making the present results directly applicable to experimentally realizable setups. On the other hand, low-energy pulses in the context of LWFA require ultrashort durations, for which carrier–envelope phase effects can influence the wakefield dynamics and electron acceleration [57, 58, 59, 60, 61]. The impact of these effects, as well as the corresponding scalability of the optimization to high-energy, long-pulse regimes, will be examined in future work.
@@ 5 Particle-in-cell simulations
To isolate the acceleration dynamics from the specifics of various injection mechanisms, an external beam of test electrons is introduced. The test electrons move self-consistently according to their charge and mass but do not contribute any current to the plasma, effectively acting as passive tracers. This approach is valid for regimes well below the beam-loading threshold [62, 63], where the influence of the accelerated electrons on the wakefield structure is negligible. The initial velocity of the test electrons is set to $v_{0}=c\sqrt{1-\omega_{\mathrm{p}}^{2}/\omega_{0}^{2}}$, approximately matching the phase velocity of the plasma wave to ensure efficient electron trapping. The beam is initialized as a cylinder, with length spanning the entire simulation domain and radius set to $2w_{0}$, thereby providing full overlap with the generated wakefield structure. Consequently, a continuous range of electron injection phases within the wakefield is sampled simultaneously in a single simulation. It is important to emphasize that the results obtained in the present manuscript remain valid for any experimentally relevant electron injection mechanism, provided that the approximate injection location and the injected charge can be controlled to a reasonable extent; an example of such a mechanism is nanoparticle-assisted injection [64, 65], which was employed in our previous study on the electron energy optimization [32].
@@ 6 Bayesian optimization
In our setup, a Gaussian process (GP) [71] is employed as the surrogate model, following standard practice in BO. A GP is defined as a collection of random variables for which any finite subset have a jointly Gaussian distribution. Formally, a function $f(\mathbf{x})$ drawn from a GP is denoted as
$f\left(\mathbf{x}\right)\sim\mathcal{GP}\left(m(\mathbf{x}),k(\mathbf{x},\mathbf{x}^{\prime})\right),$ (17)
where $m(\mathbf{x})$ is the mean function and $k(\mathbf{x},\mathbf{x}^{\prime})$ is the covariance function (or kernel). The mean function represents the expected value of the function at the input $\mathbf{x}$,
$m\left(\mathbf{x}\right)=\mathbb{E}\left[f\left(\mathbf{x}\right)\right].$ (18)
The kernel specifies the correlation between function values evaluated at inputs $\mathbf{x}$ and $\mathbf{x}^{\prime}$,
$k\left(\mathbf{x},\mathbf{x}^{\prime}\right)=\mathbb{E}\left[\left(f\left(\mathbf{x}\right)-m\left(\mathbf{x}\right)\right)\left(f\left(\mathbf{x}^{\prime}\right)-m\left(\mathbf{x}^{\prime}\right)\right)\right].$ (19)
The choice of kernel reflects prior assumptions regarding the smoothness and overall behavior of the objective function. In this work, we adopt the radial basis function (RBF) kernel [72],
$k\left(\mathbf{x},\mathbf{x}^{\prime}\right)=\exp{\left(-\frac{1}{2}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)^{\top}\Theta^{-2}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)\right)},$ (20)
where $\Theta$ denotes the length-scale hyperparameter governing the rate at which correlations decay with distance in the input space. The value of $\Theta$ is determined by maximizing the log marginal likelihood of the GP. The RBF kernel assumes the objective function to be infinitely differentiable—an appropriate choice given the expected smooth dependence of the output on the input parameters.
The acquisition function employed is the Monte Carlo–based batched upper confidence bound ($q$-UCB) [73, 74]. The $q$-UCB acquisition function includes a tunable hyperparameter $\beta_{q\text{-}\mathrm{UCB}}$, which governs the balance between exploration and exploitation. We set $\beta_{q\text{-}\mathrm{UCB}}=10$ and keep it constant throughout the optimization process, as this value has proven to provide a good balance in a series of preliminary tests (see Appendix A for details). Furthermore, a batched acquisition strategy allows multiple trials to be selected and evaluated concurrently.
Simulations considered within the present work yield noise-free observations. When extending the proposed approach to situations in which the objective function exhibits stochastic behavior, as occurs in experimental settings (e.g., when employing colliding-pulse injection [75], where counter-propagating laser pulses induce stochastic electron heating), the Bayesian optimization framework must be adapted to explicitly account for observation noise.
@@
Based on our previous work [32], the optimum is expected to lie within these intervals; hence, the parameter limits are not dynamically adjusted during the optimization process. The single optimization objective is the maximum test-electron macroparticle energy achieved in the simulation. Additionally, the acceleration length—defined as the propagation distance in plasma at which the maximum energy is reached—is recorded for post-analysis, although it is not included in the optimization objective. We do not attempt to distinguish among the individual mechanisms that determine the acceleration length (e.g., electron dephasing, pump depletion), as these processes are typically closely interconnected.
@@ 7 Optimization results
This corresponds to the following set of initial laser and plasma parameters: $a_{0}\approx 2.3$, $\tau_{0}\approx 9.5~\mathrm{fs}$ (yielding $\mathcal{P}_{0}\approx 1~\mathrm{TW}$), $w_{0}\approx 3~\mathrm{\upmu m}$, and $n_{\mathrm{e}}\approx 3\times 10^{19}~\mathrm{cm^{-3}}$. Across the explored parameter space, the maximum electron energy ranges from $\approx 25$ to $77~\mathrm{MeV}$, with the corresponding acceleration distance varying between $\approx 19$ and $192~\mathrm{\upmu m}$. For additional information on the uncertainty quantification of the GP models and their comparison with classical analytical models, see Appendices B and C, respectively.
@@
This indicates that near-maximum electron energies can be produced across several different sets of initial parameters. For instance, in the present $10~\mathrm{mJ}$ case, electrons in the top $5\%$ of the energy distribution are obtained either for $\tau_{0}\approx 9.8~\mathrm{fs}$, $w_{0}\approx 2.4~\mathrm{\upmu m}$, and $n_{\mathrm{e}}\approx 3.4\times 10^{19}~\mathrm{cm^{-3}}$, or for $\tau_{0}\approx 8.1~\mathrm{fs}$, $w_{0}\approx 3.3~\mathrm{\upmu m}$, and $n_{\mathrm{e}}\approx 2.9\times 10^{19}~\mathrm{cm^{-3}}$. Furthermore, by fixing the pulse duration and plasma density near their optimal values (i.e., $\tau_{0}\approx 9.5~\mathrm{fs}$ and $n_{\mathrm{e}}\approx 3\times 10^{19}~\mathrm{cm^{-3}}$), variations of the focal spot from $\approx 2.4$ to $3.3~\mathrm{\upmu m}$ still yield near-maximum electron energies. These findings demonstrate that operation within the optimal LWFA regime can be (i) realized through multiple parameter combinations, providing substantial flexibility for experimental implementation, and (ii) maintained without the need for precisely tuned laser or plasma parameters, thereby significantly relaxing the operational constraints.
@@
The simulation results indicate that achieving efficient self-guiding under realistic conditions requires accounting for the self-consistent evolution of the laser pulse and the plasma. As discussed in Ref. 32, this implies that (i) the initial laser power should be set slightly above the critical power and (ii) the pulse duration be somewhat longer than the resonant duration for linear plasma wave excitation. We note that, for a Gaussian pulse with $a_{0}\ll 1$, the resonant duration can be calculated analytically as $\tau_{r}\omega_{\mathrm{p}}\approx 2\sqrt{2\ln{2}}\approx 2.35$ [78].
In this work, we build upon our previous study [32], where BO was applied to the case of $\kappa=2$, the electron injection was realized using the nanoparticle-assisted method [64, 65], and all other laser and plasma parameters were kept identical. The adoption of advanced numerical techniques in the present analysis enabled a more thorough exploration of the parameter space, resulting in a maximum electron energy almost 15% higher than reported in Ref. 32.
@@ 8 Conclusion
The optimization revealed that a self-guided LWFA driven by a $10~\mathrm{mJ}$ laser pulse in a uniform plasma is, in principle, capable of producing electrons with energies approaching $80~\mathrm{MeV}$ over an acceleration distance below $200~\mathrm{\upmu m}$. The maximum energy was achieved for a proportionality parameter close to the widely accepted factor of 22 ($\pm 0.4$). The underlying physics is straightforward: efficient self-guiding requires a balance between the plasma restoring force and the laser ponderomotive force. Furthermore, electrons with energies near the maximum value were obtained across a relatively broad range of input parameters and without the need for their precise tuning, i.e., the balance between these forces need not be exact because LWFA is a dynamic, self-adjusting process. For instance, when the pulse duration and plasma density are set to $9.5~\mathrm{fs}$ and $3\times 10^{19}~\mathrm{cm^{-3}}$, respectively, varying the focal spot between $2.4$ and $3.3~\mathrm{\upmu m}$ still yields near-maximum electron energies. This provides substantial flexibility for experimental implementation and significantly relaxes the operational constraints associated with self-guided LWFA. Although the optimization is demonstrated for a relatively low-energy driver pulse, the parameter space is defined in terms of dimensionless quantities, for which the location and extent of the optimal operating region are expected to depend only weakly on the laser energy due to the self-similar nature of the underlying physics.
@@
[…]
The data that support the findings of this study are openly available at the following URL/DOI: https://doi.org/10.5281/zenodo.18610970. [79]
Appendix A. Acquisition function hyperparameter tuning
The hyperparameter $\beta_{q,\mathrm{UCB}}$ of the $q$-UCB acquisition function controls the trade-off between exploration and exploitation in the Bayesian optimization procedure, and there is no universally optimal choice for its value. Instead, the appropriate setting depends on factors such as the available evaluation budget and the specific optimization objective. For example, in scenarios with a very limited number of evaluations (e.g., costly experiments or real-time tuning), it is often preferable to emphasize exploitation in order to identify a satisfactory solution rapidly, even at the risk of missing the global optimum. Conversely, in cases where robustness or the global structure of the objective landscape is of interest, stronger exploration may be favored.
In the present study, we employ a relatively large evaluation budget of 500 PIC simulations, which allows for a balanced strategy combining both exploration and exploitation. Based on preliminary tests, we found $\beta_{q,\mathrm{UCB}}=10$ to provide a suitable compromise under these conditions. To support this choice, we include additional results showing optimization runs with $\beta_{q,\mathrm{UCB}}=1$ (Fig. 4) and $\beta_{q,\mathrm{UCB}}=100$ (Fig. 5), using an evaluation budget of 300 PIC simulations. These results demonstrate that smaller values of $\beta_{q,\mathrm{UCB}}$ lead to rapid convergence toward local optima with limited exploration of the parameter space, whereas larger values overly emphasize exploration at the expense of convergence. The chosen value $\beta_{q,\mathrm{UCB}}=10$ therefore represents a practical and well-balanced choice for the optimization problem considered here.
Figure 4: (a) Locations of the trials in the parameter space and the corresponding maximum electron energies, $\mathcal{E}_{\mathrm{e,max}}$, and (b) maximum electron energy obtained in each evaluation (blue) and the evolution of the cumulative maximum (red) for $\beta_{q,\mathrm{UCB}}=1$.
Figure 5: (a) Locations of the trials in the parameter space and the corresponding maximum electron energies, $\mathcal{E}_{\mathrm{e,max}}$, and (b) maximum electron energy obtained in each evaluation (blue) and the evolution of the cumulative maximum (red) for $\beta_{q,\mathrm{UCB}}=100$.
Appendix B. Uncertainty quantification of surrogate models
Figure 6: Standard deviation of the GP model for the maximum electron energy, $\sigma(\mathcal{E}_{\mathrm{e,max}})$, and the corresponding acceleration distance, $\sigma(l_{\mathrm{acc}})$, based on the results of 500 PIC simulations for three different values of proportionality parameter $\kappa$, (a) $\kappa=1.5$, (b) $\kappa=2.0$, and (c) $\kappa=2.5$.
GP models quantify predictive uncertainty through the posterior standard deviation, which provides a fundamental measure of confidence in the model predictions. A high standard deviation indicates large uncertainty, typically occurring in regions far from the training data, whereas a low standard deviation corresponds to high predictive confidence near sampled points.
To illustrate the spatial distribution of predictive uncertainty, Fig. 6 shows the standard deviation of the GP model for the maximum electron energy, $\sigma(\mathcal{E}_{\mathrm{e,max}})$, and the corresponding acceleration distance, $\sigma(l_{\mathrm{acc}})$ for $\kappa=1.5$ [panel (a)], $\kappa=2.0$ [panel (b)], and $\kappa=2.5$ [panel (c)]. Within the parameter space defined by Eq. (21), the maximum values of $\sigma(\mathcal{E}_{\mathrm{e,max}})$ and $\sigma(l_{\mathrm{acc}})$ are $\approx 2.9\,\mathrm{MeV}$ and $11.9\,\mathrm{\upmu m}$, respectively. In the optimal LWFA operating regime defined by Eq. (23), the standard deviation of the predicted maximum electron energy remains below $1\,\mathrm{MeV}$, indicating that the excellent performance predicted in this region is robust and supported by high model confidence.
Appendix C. Comparison with classical analytical models
To place the results obtained in this work within the context of current physical understanding, we compare the GP surrogate models for the maximum electron energy and the corresponding acceleration distance with the classical, well-established analytical expressions presented in Ref. 20. In that work, the maximum electron energy and the corresponding acceleration distance are given by
$\frac{\mathcal{E}_{\mathrm{e,max}}^{\prime}}{m_{\mathrm{e}}c^{2}}\approx\frac{2}{3}\left(\frac{n_{\mathrm{e}}}{n_{\mathrm{cr}}}\right)^{-1}a_{0}$ (24)
and
$k_{0}l_{\mathrm{acc}}^{\prime}\approx\frac{4}{3}\left(\frac{n_{\mathrm{e}}}{n_{\mathrm{cr}}}\right)^{-3/2}\sqrt{a_{0}},$ (25)
respectively. Since Eqs. (24) and (25) are derived in the matched regime, we use Eqs. (14) and (16) (with $\kappa=2$) to express them, without loss of generality, in terms of the dimensionless parameters $\mathcal{P}_{0}/\mathcal{P}_{\mathrm{cr}}$ and $\tau_{0}\omega_{\mathrm{p}}$:
$\frac{\mathcal{E}_{\mathrm{e,max}}^{\prime}}{m_{\mathrm{e}}c^{2}}\approx\frac{4}{3}\left(\frac{2\sqrt{\pi\ln{2}}}{\tau_{0}\omega_{\mathrm{p}}}\right)^{2/3}\left(\frac{\mathcal{P}_{0}}{\mathcal{P}_{\mathrm{cr}}}\right)^{-1/3}\left(\frac{\mathcal{E}_{0}}{\overline{\mathcal{E}}}\right)^{2/3},$ (26)
$k_{0}l_{\mathrm{acc}}^{\prime}\approx\frac{4\sqrt{2}}{3}\left(\frac{2\sqrt{\pi\ln{2}}}{\tau_{0}\omega_{\mathrm{p}}}\right)\left(\frac{\mathcal{P}_{0}}{\mathcal{P}_{\mathrm{cr}}}\right)^{-5/6}\left(\frac{\mathcal{E}_{0}}{\overline{\mathcal{E}}}\right).$ (27)
Equations (26) and (27) are plotted in Fig. 7, together with the regions in which they agree with the GP model shown in panel (b) of Fig. 1 to within a 20% relative error. It can be seen that Eqs. (24) and (25) are well applicable along narrow stripes in the parameter space, corresponding to values of $\mathcal{P}_{0}/\mathcal{P}_{\mathrm{cr}}$ and $\tau_{0}\omega_{\mathrm{p}}$ between approximately 1 and 2. However, these simplified analytical models cannot capture the full nonlinear structure of the objective landscape and therefore fail to identify the optimal LWFA operating regime revealed by the present work.
Figure 7: Analytical model predictions for the maximum electron energy, $\mathcal{E}_{\mathrm{e,max}}^{\prime}$, and the corresponding acceleration distance, $l_{\mathrm{acc}}^{\prime}$, as presented in Ref. 20. The green regions indicate parameter-space areas where the analytical model agrees with the GP surrogate model shown in panel (b) of Fig. 1 to within a 20% relative error.